One long wire carries current 18.0 A to the left along the x axis. A second long wire carries current 78.0 A to the right along the line (y = 0.280 m, z = 0).

(a) Where in the plane of the two wires is the total magnetic field equal to zero?One long wire carries current 18.0 A to the left along the x axis. A second long wire carries current 78.0 A to the right alo
Answer #1

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..

A) Y = -0.084 m

б. 2 Ә(So |-|о 18А = = Mo 4 (8 Мь 18 2т ( 0 - 29 +y) 78_ 18 21 (1) р. 29y р, орчь»

B)

F = qVBnet

= 2*10^-6* 150 * ( 4pi*10^-7 *78 / ( 2*pi*0.28) + 4pi*10^-7 *18 / ( 2*pi*0.10) )

= 3.68*10-8 N answer

C)

qE = 3.68*10-8

E = 3.68*10^-8/ ( 2*10^-6)

= 0.0184 N/C answer